OpenStudy (anonymous):

I need helps on a question! Simplify the radical expression. Its the square root of 20.

OpenStudy (anonymous):

\(20=4\times 5\) so \[\sqrt{20}=\sqrt{4\times 5}=\sqrt{4}\times \sqrt{5}=2\sqrt{5}\]

OpenStudy (anonymous):

Omg thanks so much!

OpenStudy (anonymous):

yw